### v1

# Themis: Towards Flexible and Interpretable NLG Evaluation

Abstract
The evaluation of natural language generation (NLG) tasks is a significant and longstanding research issue. With the recent emergence of powerful large language models (LLMs), some studies have turned to LLM-based automatic evaluation methods, which demonstrate great potential to become a new evaluation paradigm following traditional string-based and model-based metrics. However, despite the improved performance of existing methods, they still possess some deficiencies, such as dependency on references and limited evaluation flexibility. Therefore, in this paper, we meticulously construct a large-scale NLG evaluation corpus NLG-Eval with human and GPT-4 annotations to alleviate the lack of relevant data in this field. Furthermore, we propose Themis, an LLM dedicated to NLG evaluation, which has been trained with our designed multi-perspective consistency and rating-oriented preference alignment methods. Themis can conduct flexible and interpretable evaluations without references, and it exhibits superior evaluation performance on various NLG tasks, simultaneously generalizing well to unseen tasks and surpassing other evaluation models, including GPT-4.
Themis: Towards Flexible and Interpretable NLG Evaluation
Xinyu Hu, Li Lin, Mingqi Gao, Xunjian Yin, Xiaojun Wan Wangxuan Institute of Computer Technology, Peking University {huxinyu,gaomingqi,xjyin,wanxiaojun}@pku.edu.cn efsotr_l@stu.pku.edu.cn
1 Introduction
Automated evaluation is crucial for natural language generation tasks, as it measures the performance of related models and consequently promotes the development of NLG research. In the early years, traditional string-based evaluation metrics, such as BLEU (Papineni et al., 2002), were commonly used. Despite their convenience, surface-level matching cannot reliably evaluate texts as they are easily affected by perturbations (He et al., 2023), and previous work (Sulem et al., 2018a) has indicated their low correlation with human evaluations. With the development of pre-trained language models and related corpora, pre-trained model-based evaluation metrics such as BARTScore (Yuan et al., 2021) and COMET (Rei et al., 2020) have then been proposed. But their performance remains unsatisfactory compared with human evaluation, and they cannot conduct evaluations on customized aspects, like usefulness.
Figure 1: Themis is capable of evaluating various NLG tasks according to flexible evaluation aspects and criteria without references and providing corresponding analyses along with the evaluation ratings.
Recently, large language models (LLMs) such as ChatGPT and LLaMA (Touvron et al., 2023) have emerged and demonstrated unprecedented performance in instruction following and open-domain generation, which shows great potential for LLM-based automated evaluation to become a new paradigm (Wang et al., 2023a; Gao et al., 2024). Existing related studies can be divided into two major categories: directly prompting LLMs with different optimization methods (Liu et al., 2023b; Chiang and Lee, 2023b; Liu et al., 2024a; Kocmi and Federmann, 2023b) and fine-tuning LLMs with annotated data (Li et al., 2023a; Jiang et al., 2023; Xu et al., 2023; Kim et al., 2023). However, these approaches still have some limitations: the first category often relies on proprietary LLMs such as GPT-4, which are high-cost and possibly irreproducible, while the second category tends to have weaknesses of dependence on references, lacking flexibility, or poor interpretability.
In this paper, we propose Themis, an 8B-parameter LLM specifically designed and trained for NLG evaluation with more comprehensive capabilities. Our Themis can evaluate various NLG tasks, including uncommon ones like question-answering evaluation (Versatility), in a reference-free manner (Independence). Moreover, it allows for specific and customized evaluation aspects and criteria, including overall quality and more fine-grained aspects (Flexibility), and its evaluation contains corresponding analysis and explanation together with the rating (Interpretability), as shown in Figure 1. We believe that an ideal evaluator should be convenient to use and possess these characteristics. The comparison between related methods and Themis is shown in Table 1.
To obtain high-quality training data, we extensively surveyed existing studies on NLG, collecting and filtering out 58 evaluation datasets with human annotations across 9 common NLG tasks. Given the importance of evaluation criteria, we meticulously proofread each dataset and manually supplemented the missing descriptions. In addition, GPT-4 is treated as an additional annotator to supplement and validate the collected data while providing evaluation analyses. Ultimately, we constructed a large-scale evaluation corpus, NLG-Eval, which contains about 0.5 million samples. These resources are released to alleviate the data scarcity in the field and facilitate relevant research.
Furthermore, we proposed a multi-perspective consistency verification method to select relatively more reliable data from the constructed NLG-Eval corpus. Empirical methods are also employed to increase the distribution balance and diversity of data, resulting in approximately 67K training samples. Moreover, we designed specific preference alignment, guiding the construction and utilization of preference data through evaluation ratings, to improve the evaluation capabilities of the fine-tuned model. Experimental results show that our Themis achieves better overall evaluation performance over previous evaluation models on common NLG tasks, including summarization, story generation, data-to-text, and so on. We also conducted more in-depth analyses, including generalization tests on unseen tasks like the instruction-following evaluation as well as aspect-targeted perturbation tests to verify the reliability of our model.
Method Vers. Inde. Flex. Inte. Open.
UniEval
G-Eval
X-Eval
Prometheus
Auto-J
InstructScore
TIGERScore
Themis (Ours)
Table 1: Comparisons of our Themis with currently common evaluation models, including UniEval (Zhong et al., 2022), G-Eval (Liu et al., 2023b), X-Eval (Liu et al., 2023a), Prometheus (Kim et al., 2023), Auto-J (Li et al., 2023a), InstructScore (Xu et al., 2023) and TIGERScore (Jiang et al., 2023). Vers., Inde., Flex., Inte., and Open. represent versatility, independence, flexibility, interpretability, and open-source, respectively.
Overall, our main contributions are as follows:
  • We construct a large-scale NLG evaluation corpus, including about 0.5 million samples and 58 datasets across 9 NLG tasks, with detailed meta-information, aspect criteria, and evaluations.
  • We propose Themis, an LLM dedicated to NLG evaluation, which has been trained through our specific consistency and alignment methods and possesses versatility, independence, flexibility, and interpretability.
  • Extensive experiments demonstrate the superior evaluation performance of Themis in common NLG tasks, as well as good generalization and robustness. Our model and relevant resource are released in https://github.com/PKU-ONELab/Themis to facilitate related research.
2 NLG-Eval Corpus
Despite abundant data on NLG tasks, the corresponding high-quality evaluation data remains scarce and scattered due to the high cost of professional human annotations. Previous related methods either used a small amount of human annotations to train regression models, such as UniEval (Zhong et al., 2022) and X-Eval (Liu et al., 2023a), resulting in limited task coverage and a lack of evaluation analyses, or they entirely relied on LLMs to generate synthetic data, which raised reliability concerns (Hu et al., 2024). To address this challenge, we formally defined the NLG evaluation task and clarified the involved elements. Subsequently, we surveyed a large number of existing related studies and compiled 58 evaluation datasets with human annotations across 9 NLG tasks, totaling 0.5 million samples. They have undergone our meticulous proofreading, and the missing but critical content has been supplemented, such as evaluation criteria. Additionally, we utilize the competent GPT-4 for supplementary evaluations, including analyses and ratings. The corpus has also been equipped with much meta-information, aiming to promote the development of related research.
Figure 2: The NLG evaluation tasks and corresponding evaluation aspects in our NLG-Eval corpus.
2.1 Definitions
NLG tasks typically require language models to generate the corresponding output $t$ based on the given input $i$ and task requirements. The input in specific tasks may consist of multiple components; for example, in controlled generation, it is composed of the generation requirements and the constraint labels. NLG evaluation involves evaluating $t$ in conjunction with $i$ through the evaluation model or method $E$, using evaluation criteria $c$ and references $r$ if available. When they are absent, the evaluation defaults to general and reference-free evaluation, respectively. Evaluation results come in various forms, primarily focusing on rating $s$ and additional information $a$ such as analyses. The formalization of the evaluation process is as follows, with square brackets indicating optional operations:
$s,[a]=E(t,i,[r],[c])$
In the corpus construction, we collect and annotate the aforementioned elements for each sample.
2.2 Data Collection
To improve the coverage and diversity of the corpus, we conducted a comprehensive survey of related work on NLG tasks, including summarization, dialogue response generation, data-to-text, and so on. We collected datasets with human evaluation ratings and performed meticulous manual checking and filtering, ultimately obtaining 58 high-quality datasets across 9 different NLG tasks, which contain approximately 0.5 million samples. However, we discovered that many datasets, despite having evaluation aspects, lacked specific criteria or definitions, which could lead to confusion and ambiguity in evaluations. Such concerns have also been pointed out by Zhou et al. (2022); Hu et al. (2024). So we analyzed the involved datasets, taking samples, human labels, and information from related papers into account, and manually supplemented the criteria.
Although human evaluation is often considered the most reliable, many datasets have sparse or inconsistent human annotations, leading to questions about their accuracy. Therefore, we utilized GPT-4 to cross-verify with human evaluations to improve reliability, as it has been shown to perform well on evaluation (Zheng et al., 2023). On the other hand, it can provide evaluation analyses beyond just ratings and increase interpretability, which is lacking in human evaluation. More implementation details can be found in Appendix A. Additionally, our subsequent training did not use references because many datasets lack references or only have automatically generated references that are unsatisfactory (Pu et al., 2023). And the references are also difficult to obtain in practical evaluations. The tasks included in our NLG-Eval corpus and their respective evaluation aspects are shown in Figure 2, with more statistics and detailed information presented in Appendix B.
3 Methodology
Existing studies have highlighted that data quality is more crucial than quantity in LLM fine-tuning (Li et al., 2023c; Zhou et al., 2023). Therefore, we adopted empirical methods to improve data diversity and distribution balance and designed a multi-perspective consistency verification approach to obtain about 67K high-quality samples from the constructed corpus. Then they are used for supervised fine-tuning based on open-source LLMs. To further enhance the evaluation capability and alignment of the model, we proposed a preference data construction and training method guided by evaluation ratings to improve the fine-tuned model.
3.1 Diversity and Balance
Similar to many studies on LLM training, we initially employed empirical approaches to sample data from the original corpus, ensuring adequate diversity and distribution balance, which is generally considered beneficial for model training and generalization. To be specific, we treated the NLG task, evaluation aspect, and rating triplets as identifiers of data categories to cover as comprehensive a distribution as possible. And we sampled 100 pieces of data from each category while also considering the semantic diversity of the texts involved. Given the insufficiency of data in some categories and the requirement to introduce diversity in evaluation analyses and aspect criteria, we conducted multiple samplings of analyses differently for each category to further balance the data distribution, simultaneously rephrasing the evaluation aspect criteria. In the end, we obtained a dataset of approximately 67K for supervised fine-tuning. The detailed sampling method for both samples and evaluation analyses will be introduced in the next subsection.
3.2 Multi-perspective Consistency Verification
During the sampling process, most categories actually had sufficient samples, and each sample also involved multiple annotated evaluations. To take full advantage of such sufficient data resources and enhance the reliability of the training data, we verified the consistency of each sample and their evaluations from three different perspectives, ultimately filtering out the potentially high-quality data.
Self-Consistency
When dealing with complex tasks like mathematical problems and logical reasoning, LLMs often employ the chain of thought (CoT) (Wei et al., 2022) method. It involves generating intermediate steps before providing the final result, which generally leads to higher-quality responses. Furthermore, by performing multiple diverse samplings and selecting the most frequently occurring final result from the responses, the model’s performance can be further improved (Wang et al., 2023c). This is referred to as self-consistency, which can be calculated as the proportion of the most frequently occurring result to the total number of samplings. A higher self-consistency indicates greater certainty of the model itself and, thus, higher reliability of the corresponding responses. In our evaluation annotation, GPT-4 was required to generate an analysis before assigning a rating, similar to a CoT process, which allows using self-consistency for data filtering. So we first collected samples where at least half of the evaluation ratings are consistent, and such ratings are named consistent ratings. Our training data would be selected from these samples and their evaluations that include consistent ratings.
Cross-Validation
The evaluation ratings serve as the most critical supervision signal in evaluation, making their accuracy paramount. Unlike previous related work that relied entirely on human labels or LLM-generated labels, we comprehensively considered the evaluations from both sources to conduct cross-validation. As for each sample, the corresponding consistent rating is regarded as its rating from GPT-4 and scaled to the same range as the human rating. We prioritized samples where the two evaluation ratings were close, which reflected the high consistency between evaluations from humans and GPT-4 and also indicated the potential strong reliability of the samples.
Evaluation Inspection
Hu et al. (2024) indicated that although current LLMs possess exciting evaluation capabilities, they may encounter issues of confusing aspect criteria, affecting the accuracy of their analyses and ratings. To address this, we propose two specific criteria for inspecting the evaluations themselves: consistency between the evaluation analysis and rating, and consistency between the evaluation analysis and aspect. We designed additional instructions to prompt GPT-4 to re-evaluate the candidate evaluations, whose details could be found in Appendix A, and prioritized those performing well from both perspectives.
3.3 Preference Alignment
With the rise of InstructGPT (Ouyang et al., 2022), reinforcement learning from human feedback (RLHF), as one of the key technologies, has garnered wide attention and been applied in many subsequent LLMs with great success. So we have specifically modified DPO (Rafailov et al., 2023), a commonly-used implementation of RLHF, based on our evaluation scenarios, using evaluation ratings to guide the construction and training of preference data. The preference alignment was further conducted after the supervised fine-tuning of the model.
3.3.1 Preference Construction
Although the DPO method does not require training a separate reward model, it still necessitates preference pairs to convey preference information. Benefiting from multiple evaluations of each sample, we can directly regard evaluations that meet the consistent rating introduced in the Section 3.2 as the chosen response and those that do not as the rejected response to construct preference pairs. The quality gap between the two evaluations can be reflected by the difference in their evaluation ratings, as consistency ratings are deemed more reliable. And our preference data is constructed on the previously used training data without involving additional samples.
3.3.2 Rating-guided DPO
According to DPO, we can reparameterize the reward function $r$ using the policy as follows:
$r(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}+\beta\log Z(x)$
where $\pi_{\theta}$ is the policy model, $\pi_{\mathrm{ref}}$ is the reference model, typically a model undergone supervised fine-tuning, $\beta$ is the parameter controlling the degree of deviation between the policy and the reference, and $Z(x)$ is the partition function. The Bradley-Terry (BT) (Bradley and Terry, 1952) model is employed to model preferences, yielding:
$p(y_{1}\succ y_{2}|x)=\sigma(r(x,y_{1})-r(x,y_{2}))$
For a pair of evaluation responses $(y_{1},y_{2})$, if we prefer $y_{1}$ to $y_{2}$, then the larger the evaluation rating difference between $y_{1}$ and $y_{2}$, the greater the preference difference $r(x,y_{1})-r(x,y_{2})$ should be. Therefore, to treat each preference pair more equally, we modify the BT model by subtracting a value proportional to the rating difference between $y_{1}$ and $y_{2}$ from $r(x,y_{1})-r(x,y_{2})$, thereby compensating for the prior preference difference caused by the rating difference, and then obtain:
$p^{*}(y_{1}\succ y_{2}|x)=\sigma(r(x,y_{1})-r(x,y_{2})-\alpha|R(y_{1})-R(y_{2})|)$
where $R(y)$ denotes the evaluation rating in $y$. Finally, the maximum likelihood of new preference model serves as our optimization objective:
$\mathcal{L}^{*}_{\mathrm{DPO}}(\pi_{\theta};\pi_{\mathrm{ref}})\hskip 30.0pt$
$=-\mathbb{E}_{(x,y_{c},y_{r})\sim\mathcal{D}}\big{[}\log\sigma\big{(}$ $\beta\log\frac{\pi_{\theta}(y_{c}|x)}{\pi_{\mathrm{ref}}(y_{c}|x)}$
$-\beta\log\frac{\pi_{\theta}(y_{r}|x)}{\pi_{\mathrm{ref}}(y_{r}|x)}-$ $\alpha|R(y_{c})-R(y_{r})|\big{)}\big{]}$
where $(x,y_{c},y_{r})$ contains the input, the chosen response, and the rejected response, respectively.
4 Experiments
4.1 Benchmarks
We conducted extensive experiments to evaluate our Themis across six commonly used datasets for different NLG tasks, including SummEval (Fabbri et al., 2021) for summarization, Topical-Chat (Gopalakrishnan et al., 2019) for dialogue response generation, SFRES&SFHOT (Wen et al., 2015) for data-to-text, QAGS (Wang et al., 2020) for factuality, MANS (Guan et al., 2021) for story generation, and WMT23 (Freitag et al., 2023) for machine translation. These datasets were collected into our NLG-Eval corpus but were excluded during the construction of our training data to ensure fairness.
Method SummEval Topical-Chat SFHOT&RES QAGS MANS WMT23 Average
$\rho$ $\tau$ $r$ $\rho$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $r$ $\rho$ $\rho$
Traditional Metrics
BLEU 0.075 0.057 0.356 0.388 0.024 0.018 - - 0.032 0.009 -0.130 0.021 -
ROUGE 0.152 0.120 0.393 0.412 0.101 0.076 - - -0.002 0.156 0.081 0.151 -
BARTScore 0.329 0.261 0.067 0.086 0.208 0.156 0.425 0.347 0.350 0.260 0.091 0.118 0.253
BERTScore 0.231 0.182 0.388 0.394 0.139 0.105 - - 0.285 0.163 0.123 0.219 -
BLEURT 0.152 0.118 0.384 0.388 0.244 0.184 - - 0.138 0.221 0.163 0.263 -
CometKiwi 0.228 0.180 0.353 0.340 0.251 0.186 0.094 0.074 0.251 0.176 0.413 0.343 0.251
UniEval 0.474 0.377 0.533 0.577 0.282 0.211 - - - - - - -
Prompting LLM
G-Eval (GPT-3.5) 0.409 0.323 0.574 0.585 - - 0.461 0.337 - - - - -
G-Eval (GPT-4) 0.523 0.423 0.575 0.588 - - 0.611 0.532 - - - - -
GPT-3.5 0.416 0.340 0.592 0.578 0.306 0.239 0.431 0.356 0.328 0.295 0.388 0.347 0.401
GPT-4 0.511 0.423 0.770 0.746 0.320 0.260 0.637 0.532 0.473 0.260 0.496 0.437 0.521
Fine-tuned LLM
X-Eval 0.480 0.362 0.539 0.605 0.303 - 0.578 - - - - - -
Prometheus-13B 0.163 0.142 0.435 0.434 0.173 0.142 - - 0.007 0.146 0.144 0.129 -
Auto-J-13B 0.198 0.172 0.427 0.425 0.141 0.120 0.226 0.209 0.380 0.284 0.128 0.104 0.246
TIGERScore-13B 0.384 0.334 0.334 0.346 0.200 0.175 0.504 0.446 0.231 0.207 0.277 0.248 0.319
InstructScore-7B 0.258 0.226 0.269 0.241 0.247 0.210 - - 0.298 0.168 0.213 0.219 -
Themis-8B (ours) 0.553 0.499 0.733 0.725 0.333 0.284 0.684 0.613 0.551 0.501 0.431 0.405 0.542
[…]
4.2 Experimental Settings
We chose Llama-3-8B (Meta, 2024) as our base model for supervised fine-tuning and preference alignment, and more training details are described in Appendix C. When assessing the evaluation capability of different models, we calculated the correlation between ratings from model and human using Pearson ($r$), Kendall ($\tau$), and Spearman ($\rho$) correlation coefficients. The specific calculations and the selection of correlation coefficients vary across different datasets, and we follow the common setups from previous work. Furthermore, while previous work often employed a temperature setting of T = 1 to perform multiple sampling and aggregate the evaluation results, which can enhance performance, we believe it is costly in practice and results in the loss of evaluation analyses. Therefore, we tested our Themis with T = 0 to obtain deterministic results, with other methods using their original settings. Given that most metrics and models do not support specific evaluation aspects, we calculate the results of different aspects using their overall scores, as in previous work.
4.3 Baselines
We experimented with representative and commonly-used evaluation metrics and models from three categories. Traditional metrics include BLEU, ROUGE (Lin, 2004), BARTScore, BERTScore (Zhang et al., 2020), BLEURT (Sellam et al., 2020), CometKiwi (Rei et al., 2022), and UniEval. Methods of direct prompting LLMs include G-Eval and the evaluations with GPT-3.5 and GPT-4, implemented following Chiang and Lee (2023b). Fine-tuned LLMs include X-Eval, Prometheus, Auto-J, TIGERScore, and InstructScore. Many of them are reference-based and therefore cannot be tested on the reference-free QAGS dataset.
4.4 Main Results
The main experimental results of our Themis compared with other baselines are shown in Table 2. We present the average correlation coefficients for each task, with the complete results demonstrated in Appendix D. Our Themis achieves the best average evaluation performance on six NLG tasks and greatly surpasses other baselines including fine-tuned larger LLMs, except GPT-4. It only outperforms Themis in dialogue response and translation evaluations, likely because GPT-4 has considerably stronger dialogue and multilingual capabilities than Llama-3-8B, which are almost impossible to improve during our training. And our Themis avoids the limitations of high costs and instability in proprietary LLMs and can be conveniently used offline. Furthermore, although models such as BLEURT, X-Eval, and Prometheus conduct evaluation based on references, they still lag behind our reference-free Themis, showing its superiority.
4.5 Ablation Study
Method Avg $\rho$ Avg $\tau$ Avg $r$
BARTScore 0.253 0.197 0.262
CometKiwi 0.251 0.195 0.272
GPT-4 0.521 0.417 0.564
TIGERScore-13B 0.319 0.280 0.333
Auto-J-13B 0.246 0.211 0.262
Themis-8B (ours) 0.542 0.486 0.569
Data Filtering
    100% raw data 0.493 0.445 0.515
    50% raw data 0.478 0.432 0.505
    67K raw data 0.476 0.428 0.507
Preference Alignment
    vanilla DPO 0.528 0.474 0.556
    without DPO 0.508 0.456 0.534
Foundational Model
    Mistral-7B 0.489 0.440 0.515
    Llama-2-7B 0.478 0.430 0.508
    Llama-2-13B 0.500 0.449 0.533
Table 3: The results with different ablation settings.
We conducted ablation studies on our two main methods: consistency-based data selection for supervised fine-tuning and rating-guided preference alignment. Additionally, considering the superior performance of Llama 3 among LLMs of the same size, we experimented with other LLMs as the foundational model to verify the effectiveness of our approach. The average results for the six task datasets under different settings are shown in Table 3.
Data Selection We experimented with approximately 67K unfiltered data that has the same scale as the constructed data for supervised fine-tuning, and also trained the model using half and all of the corpus data. The results show that although more data bring some improvements, the training cost increases significantly, and there is still a substantial performance drop compared to our current model.
Preference Alignment The model trained with our rating-guided DPO method exhibits progressive improvements over the model trained with vanilla DPO, and then the model that has only been fine-tuned. It indicates that preference alignment has certain effects on the training of NLG evaluation tasks, and our specific method is better suited for evaluation scenarios.
Foundational Model Although the evaluation models trained based on other LLMs do not perform as well as our current model, the models using Llama-2 still significantly outperformed TIGERScore and Auto-J, which also use Llama-2-13B as the foundation model. It demonstrates the effectiveness of the training data and methods we introduced.
5 Reliability Analyses
To analyze and investigate the reliability of our Themis, we consider two aspects: generalizability, which means the model can effectively evaluate unseen NLG tasks during training, and robustness, which means the model can accurately evaluate texts with noise and perturbation.
Method Instruction Following Long form Question Answering Average
CLA COM COR POL REL EU FA LC OU SL
BARTScore -0.053 -0.040 0.063 -0.038 0.025 0.098 0.174 0.360 0.296 0.464 0.135
CometKiwi 0.345 0.464 0.349 0.223 0.473 0.155 0.128 0.210 0.138 0.134 0.262
GPT-3.5 0.257 0.488 0.433 0.239 0.401 0.612 0.318 0.443 0.212 0.562 0.396
GPT-4 0.257 0.653 0.573 0.338 0.455 0.767 0.431 0.237 -0.004 0.711 0.442
Auto-J 0.422 0.562 0.571 0.386 0.548 0.095 0.255 0.109 0.191 0.033 0.317
TIGERScore 0.262 0.247 0.285 0.198 0.239 -0.037 0.075 0.009 0.002 -0.217 0.106
Themis (ours) 0.414 0.381 0.685 0.349 0.500 0.835 0.538 0.700 0.365 0.684 0.545
Table 4: The results on instruction-following and long-form question-answering evaluations.
5.1 Unseen Tasks
We tested our model on two uncommon and unseen evaluation tasks: long-form question answering and instruction-following evaluation (Li et al., 2023b). The former task involves five evaluation aspects: clarity (CLA), completeness (COM), correctness (COR), politeness (POL), and relevance (REL), while the latter involves example usage (EU), factual accuracy (FA), logical coherence (LC), overall usefulness (OU), and simple language (SL). As shown in Table 4, Themis outperformed both proprietary and open-source LLMs on the overall two tasks, demonstrating good generalization. And its performance is relatively well-balanced without any particular weaknesses, unlike Auto-J, which performs well on instruction-following but poorly on long-form question-answering. Furthermore, some evaluation aspects are relatively open, which traditional reference-based metrics that do not support specific evaluation aspects cannot handle, highlighting the superiority of our Themis.
5.2 Perturbation Tests
[…]
GPT-4 1.158 0.896 0.929 0.923 0.976
Prometheus 1.095 1.433 1.051 1.485 1.266
Themis (ours) 0.841 0.748 0.698 0.860 0.787
Table 5: The results of perturbation tests.
Hu et al. (2024) has pointed out that both proprietary and open-source LLMs have certain issues with confusing evaluation aspects during NLG evaluation. Therefore, we conducted perturbation tests based on their methods to investigate the robustness of different evaluation models that support customized aspects. Specifically, we applied aspect-targeted perturbations on references in four tasks—news summarization, dialogue summarization, paraphrase generation, and data-to-text—focusing on fluency, coherence, informativeness, and consistency. The perturbations were designed to impact only the quality of the target aspect while leaving other aspects unaffected. We tested the models by comparing ratings on the perturbed texts with those on the original references, expecting little change and decreases in unaffected aspects. As shown in Table 5, Themis is closest to what is expected, with the smallest average decreases across the four tasks. On one hand, slight decreases may be caused by the fact that the designed pertubations do not entirely avoid affecting other aspects. Moreover, the evaluation of perturbed texts may indeed be challenging, necessitating further exploration to improve the evaluator.
6 Related Works
6.1 Prompting LLMs for NLG Evaluation
Previous research has evaluated text quality in various tasks by directly prompting proprietary LLMs (e.g. GPT-4) for scoring Liu et al. (2023b); Chiang and Lee (2023a); Kocmi and Federmann (2023b), comparison (Liusie et al., 2024; Wang et al., 2023d), ranking (Ji et al., 2023; Liu et al., 2023d), and error analysis (Kocmi and Federmann, 2023a; Lu et al., 2023), surpassing traditional evaluation metrics in performance and providing significant flexibility. Leiter et al. (2023) have experimented with various prompt formats and in-contextual example settings. Moreover, to better assess specific aspects, Liu et al. (2024a) have used LLMs to generate or improve the definitions of evaluation criteria for use in prompting, and Gong and Mao (2023a) have had LLMs first evaluate other relevant aspects. Additionally Wu et al. (2023); Bai et al. (2023) have enhanced evaluative capabilities through interactions and role-playing among LLMs. Although these studies are compelling, they rely on proprietary models, which are costly and pose reproducibility issues.
6.2 Fine-tuned LLM Evaluators
In response to the issues with prompting LLMs, subsequent studies have shifted to fine-tuning relatively smaller open-source large models to create specialized evaluators (Wang et al., 2023e; Xu et al., 2023; Kim et al., 2023; Wang et al., 2023b; Jiang et al., 2023; Li et al., 2023a; Ke et al., 2023; Zhu et al., 2023; Liu et al., 2023a). Their approaches are generally similar: they use existing human evaluation data or synthetic data created by proprietary LLMs, such as GPT-4, as training data to fine-tune open-source LLMs like Llama (Touvron et al., 2023). They vary in details, such as whether references are needed and whether they support customized evaluation criteria. However, without specific measures, such fine-tuned evaluators were shown to contain biases similar to those in proprietary LLMs (Hu et al., 2024).
7 Conclusion
In this paper, we propose a large-scale and comprehensive NLG evaluation corpus and an LLM dedicated to NLG evaluation, Themis. The NLG-Eval corpus encompasses 9 common NLG tasks across 58 datasets, comprising about 0.5 million samples with evaluations from both human and GPT-4. Based on the constructed corpus, we propose a specialized multi-perspective consistency verification method to filter out high-quality supervised fine-tuning data, together with preference alignment guided by evaluation ratings, to train our Themis. It can be applied to various NLG tasks with flexible and interpretable evaluation in a reference-free manner. Extensive experiments demonstrate that Themis performs well on various NLG tasks and can be generalized well to unseen tasks. Our model and resources have been released, and we hope they will promote further research in the field of NLG evaluation.
Limitations
Due to limited computational resources, we did not use larger and more powerful LLMs than Llama-3-8B as the foundation model for training, such as Llama-3-70B. Although larger models might offer stronger evaluation capabilities, their practical usage becomes inconvenient. Balancing model performance and usability requires further exploration in the future.
Acknowledgements
This work was supported by National Science Foundation of China (No. 62161160339), National Key R&D Program of China (2021YFF0901502), Ant Group Research Fund and Key Laboratory of Science, State Key Laboratory of Media Convergence Production Technology and Systems and Key Laboratory of Science, Technology and Standard in Press Industry (Key Laboratory of Intelligent Press Media Technology). Xiaojun Wan is the corresponding author.
Appendix A Settings and Prompts for Annotations
We follow the suggestions of Bsharat et al. (2023) to design the instructions and prompts for GPT-4 to annotate NLG evaluation tasks, as shown in Table 18. And similar to Chiang and Lee (2023b), we require to provide the concise and accurate analysis before the evaluation rating with a temperature setting of 1 for sampling. The prompts for re-evaluating evaluations are shown in Table 19 and Table 20, where the task descriptions are appropriately modified to prevent confusion.
Appendix B Details for NLG-Eval Corpus
Our constructed NLG-Eval corpus includes evaluation datasets on nine NLG tasks, including Controllable Generation, Data to Text, Dialogue Response Generation, Grammatical Error Correction, Machine Translation, Paraphrase Generation, Story Generation, Summarization and Text Simplification. The statistics are shown in Table 6 and more detailed meta-information is described in Table 7 to Table 17. Through such data and information, we hope to facilitate related research and advance the field of NLG evaluation.
Task #Datasets #Aspects #Samples
Controllable Generation 4 8 11299
Data to Text 6 11 36548
Dialogue Response Generation 17 17 91111
Grammatical Error Correction 3 6 41058
Machine Translation 2 6 347504
Paraphrase Generation 2 3 18299
Story Generation 6 17 12636
Summarization 10 17 61977
Text Simplification 8 8 27026
Table 6: The statistics of NLG-Eval corpus.
Dataset Size Aspect
chiang LLM Evaluation (Chiang and Lee, 2023a) 1600 Cohesiveness: How well do the sentences in the story fragment fit together?
Grammaticality: How grammatically correct is the text of the story fragment?
Likability: How enjoyable do you find the story fragment?
Relevance: How relevant is the story fragment to the story prompt?
CoEval (Li et al., 2023b) 1400 Character Development: The characters in the generated story should be well-developed.
Clarity: The generated story should be clear and easy to understand, with no confusing or ambiguous elements.
Coherence: The generated story should have a logical flow and provide closure.
Engagement: The generated story should be engaging from beginning to end.
Grammaticality: The generated story should be grammatically correct.
Length: The generated story should have an appropriate length according to the story requirement.
Relevance: The generated story should be relevant to the story requirement (beginning or topic) and the daily events it aims to capture.
Hanna (Chhun et al., 2022) 6336 Coherence: How much does the generated story make sense?
Complexity: How elaborate is the generated story, involving complex concepts, realistic characters, an intricate plot, an underlying history or circumstances, or precise descriptions?
Empathy: How well do you understand the character’s emotions in the generated story (regardless of whether you agree with them)?
Engagement: How much do you engage with the generated story?
Relevance: How well does the generated story match the story prompt?
Surprise: How surprising is the end of the generated story, with enough clues for a reasonable prediction?
Mans_roc (Guan et al., 2021) 1000 Overall Quality: The overall quality of the generated story, considering whether it has global errors like chaotic scenes (difficult to understand as a whole) and local errors, including repetitive plots (repeating similar texts), unrelated events (to the story beginning or within its own context), and conflicting logic (against common sense or with wrong causal or temporal relationship).
Mans_wp (Guan et al., 2021) 1000 Overall Quality: The overall quality of the generated story, considering whether it has global errors like chaotic scenes (difficult to understand as a whole) and local errors, including repetitive plots (repeating similar texts), unrelated events (to the story prompt or within its own context), and conflicting logic (against common sense or with wrong causal or temporal relationship).
nextchapter (Xie et al., 2023) 1300 Coherence: How well do the sentences in the generated story fit together?
Fluency: How grammatically correct is the text of the generated story?
Interestingness: How enjoyable do you find the generated story?
Logicality: How much does the generated story obey commonsense?
Relatedness: How relevant is the generated story to the story prompt?
Table 7: Datasets on Story Generation Task.
Dataset Size Aspect
CTRLEval (Ke et al., 2022) 3960 Attribute Relevance: Measure whether the generated text satisfies the attribute label.
Coherence: Measure whether the sentences in the generated text are semantically relevant to compose a coherent body, which reflects the quality of the generated text itself.
Consistency: Evaluate whether the generated text is consistent to the content prefix.
FUDGE (Yang and Klein, 2021) 2088 Fluency: Is the generated text fluent, i.e., well-written and grammatical?
PPLM (Dathathri et al., 2020) 3251 Fluency: Whether the generated text has no grammatical errors, formatting problems, or obviously ungrammatical issues (e.g., fragments, missing components) that make the text difficult to read?
InstruSum (Liu et al., 2023c) 2000 Factual Consistency: Is the summary consistent with the facts presented in the article, without contradicting or misrepresenting any information?
Irrelevant Information: Does the summary include any information that is not relevant to the summary requirement?
Missing Information: Does the summary omit any crucial information from the article concerning the summary requirement?
Overall Quality: Assess the overall quality of the summary in relation to the summary requirement.
Table 8: Datasets on Controllable Generation Task.
Dataset Size Aspect
E2E NLG (Dusek et al., 2020) 6300 Naturalness: Could the utterance have been produced by a native speaker?
Overall Quality: How is the overall quality of the utterance in terms of its grammatical correctness, fluency, adequacy and other important factors?
INLG16 (Novikova et al., 2016) 3726 Informativeness: Is this utterance informative? (i.e. do you think it provides enough useful information from the data?)
Naturalness: Is this utterance natural? (e.g. could it have been produced by a native speaker?)
Phrasing: Is this utterance well phrased? (i.e. do you like how it is expressed?)
RankMe (Novikova et al., 2018) 900 Informativeness (= adequacy): Does the utterance provide all the useful information from the meaning representation?
Naturalness (= fluency): Could the utterance have been produced by a native speaker?
Overall Quality: How do you judge the overall quality of the utterance in terms of its grammatical correctness, fluency, adequacy and other important factors?
SFRES_SFHOT (Wen et al., 2015) 6168 Informativeness: Does the utterance provide all the useful information from the meaning representation?
Naturalness: Could the utterance have been produced by a native speaker?
Overall Quality: How do you judge the overall quality of the utterance in terms of its grammatical correctness and fluency?
webnlg_2017 (Gardent et al., 2017) 5214 Fluency: Does the text sound fluent and natural?
Grammaticality: Is the text grammatical (no spelling or grammatical errors)?
Semantic Adequacy: Does the text correctly represent the meaning in the data?
webnlg_2020 (Castro Ferreira et al., 2020) 14240 Correctness: When describing predicates which are found in the data, does the text mention the correct objects and adequately introduce the subject for this specific predicate?
Data Coverage: Does the text include descriptions of all predicates presented in the data?
Fluency: Is it possible to say that the text progresses naturally, forms a coherent whole and it is easy to understand the text?
Relevance: Does the text describe only such predicates (with related subjects and objects), which are found in the data?
Text Structure: Is the text grammatical, well-structured, written in acceptable English language?
Table 9: Datasets on Data to Text task.
Dataset Size Aspect
convai2-grade (Huang et al., 2020b) 600 Coherence: The response should be coherent with the dialogue context, maintaining a good logical flow.
dailydialog-grade (Huang et al., 2020b) 300 Coherence: The response should be coherent with the dialogue context, maintaining a good logical flow.
dailydialog-gupta (Gupta et al., 2019) 500 Appropriateness: Whether the response is appropriate given the dialogue context in grammar, topic, and logic?
dailydialog-zhao (Zhao et al., 2020) 3600 Appropriateness: The response should be appropriate given the preceding dialogue.
Content Richness: The response should be informative, with long sentences including multiple entities and conceptual or emotional words.
Grammatical Correctness: The response should be free of grammatical and semantic errors.
Relevance: The response should be on-topic with the immediate dialogue history.
DialogADV (Liu et al., 2023e) 16416 Coherence: The logical and semantic coherence between the response and dialogue history (previous context).
Consistency: The logical and factual consistency between the response and dialogue history (previous context), facts also include external commonsense knowledge.
Fluency: The fluency and grammatical correctness of the response.
Relevance: The degree to response is connected or relevant to a particular topic, question, or situation of the dialogue history (previous context).
dstc10-persona_clean (Zhang et al., 2021) 19316 Appropriateness: The response should be appropriate given the preceding dialogue.
Content Richness: The response should be informative, with long sentences including multiple entities and conceptual or emotional words.
Grammatical Correctness: The response should be free of grammatical and semantic errors.
Relevance: The response should be on-topic with the immediate dialogue history.
dstc10-topical_clean (Zhang et al., 2021) 18000 Appropriateness: The response should be appropriate given the preceding dialogue.
Content Richness: The response should be informative, with long sentences including multiple entities and conceptual or emotional words.
Grammatical Correctness: The response should be free of grammatical and semantic errors.
Relevance: The response should be on-topic with the immediate dialogue history.
empathetic-grade (Huang et al., 2020b) 300 Coherence: The response should be coherent with the dialogue context, maintaining a good logical flow.
esl (Lee et al., 2020) 1242 Appropriateness: The response should be appropriate given the dialogue context, in grammar, topic, and logic.
fed-turn (Mehri and Eskénazi, 2020) 3375 Correctness: Is the response correct or was there a misunderstanding of the conversation?
Engagingness: Is the response engaging to user and fulfill the particular conversational goals implied by the user?
Fluency: Is the response fluently written and free of grammatical and semantic errors?
Interestingness: To the average person, is the response interesting?
Overall Quality: What is the overall impression of the response, and quality of and satisfaction with the conversation?
Relevance: Is the response relevant to and on-topic with the conversation history?
Semantical Appropriateness: Is the response semantically appropriate given the conversation history?
Specificity: Does the response produce unique and non-generic information that is specific to the conversation history?
Understandability: Is the response understandable?
holistic dialogue (Pang et al., 2020) 400 Coherence: Measures the meaningfulness of the response within the context of prior query.
Fluency: Measures the quality of phrasing of the response relative to a human native speaker.
humod (Merdivan et al., 2020) 19000 Fluency: The response should be written naturally and free of grammatical and semantic errors.
Relevance: The Response should be on-topic with the immediate dialogue history.
jsalt (Kong-Vega et al., 2018) 741 Appropriateness: Measure how well the response is semantically and pragmatically valid given the previous recent dialogue history.
Table 10: Datasets on Dialogue Response Generation task. (Part 1)
Dataset Size Aspect
ncm (Sedoc et al., 2019) 2461 Appropriateness: Whether the response is appropriate given the dialogue history, in grammar, topic, and logic?
persona-usr (Sedoc et al., 2019) 1800 Context Maintenance: Does the response serve as a valid continuation of the dialogue context (conversation history)?
Interestingness: Is the response dull or interesting?
Knowledge Use: Given the fact that the response is conditioned on, how well does the response use that fact?
Naturalness: Does the response seem to be something that a person would naturally say?
Overall Quality: What is the overall impression of this utterance? Please consider whether the response is understandable and natural, and how well it maintains context and uses knowledge.
Understandability: Is the response understandable given the previous dialogue context? (Not if its on topic, but for example if it uses pronouns they should make sense)
persona-zhao (Zhao et al., 2020) 900 Appropriateness: The response should be appropriate given the preceding dialogue.
topical-usr (Sedoc et al., 2019) 2160 Context Maintenance: Does the response serve as a valid continuation of the dialogue context (conversation history)?
Interestingness: Is the response dull or interesting?
Knowledge Use: Given the fact that the response is conditioned on, how well does the response use that fact?
Naturalness: Does the response seem to be something that a person would naturally say?
Overall Quality: What is the overall impression of this utterance? Please consider whether the response is understandable and natural, and how well it maintains context and uses knowledge.
Understandability: Is the response understandable given the previous dialogue context? (Not if its on topic, but for example if it uses pronouns they should make sense)
Table 11: Datasets on Dialogue Response Generation task. (Part 2)
Dataset Size Aspect
GMEG (Napoles et al., 2019) 27195 Overall Quality: Whether the corrected text is perfect, namely grammatical and not garbled?
protagolabs (Sottana et al., 2023) 1200 Grammaticality: The quality of the correction and the extent to which errors are left in the corrected text, regardless of whether they are present in the original text or they are newly introduced errors in the supposed corrected version.
Over-correction: Assess whether the correction avoids unnecessary syntax changes or being unnecessarily verbose, since there can be multiple ways to correct a text. The best correction should be done with the minimum number of edits.
Semantics: Whether the meaning of the original text is preserved following the grammatical error correction? NOTE: You should penalize corrections which change the meaning unnecessarily.
TMU-GFM (Yoshimura et al., 2020) 12663 Fluency: How natural does the corrected text sound for native speakers?
Grammaticality: The grammatical correctness of the corrected text and how comprehensible it is.
Meaning Preservation: The extent to which the meaning of the original text is preserved in the corrected text.
Table 12: Datasets on Grammatical Error Correction task.
Dataset Size Aspect
parabank (Hu et al., 2019) 11140 Fluency: Whether the paraphrase is meaningful and grammatical?
Semantic Similarity: Whether the paraphrase maintains similar semantics to the original text?
twitter para (Shen et al., 2022) 7159 Overall Quality: The paraphrase should not only maintain similar semantics to the original text, but also possess lexical or syntactic differences from the original text, with fluent and coherent content.
Table 13: Datasets on Paraphrase Generation task.
Dataset Size Aspect
WMT_zhen (Freitag et al., 2021) 346504 Accuracy: The translation should accurately represent the source text, not including information not present in the source text or missing content from the source text.
Fluency: The translation should not have incorrect punctuation and spelling, problems with grammar, internal inconsistency, or garbled characters due to incorrect encoding.
Locale Convention: The translation should not have the wrong formats for addresses, currency, dates, names, telephone numbers, and time expressions.
Overall Quality: The overall quality of the translation, including accuracy (e.g., mistranslation, omission, addition), fluency (e.g., grammar, spelling, punctuation, inconsistency), locale convention (e.g., format for names, currency, address), terminology (e.g., inappropriate or inconsistent usage), and style (e.g., stylistic problems).
Style: Does the translation have no stylistic problems?
Terminology: Whether the terminology of the translation is standard, appropriate for the context, and used consistently?
HumanMT (Kreutzer et al., 2018) 1000 Overall Quality: The overall quality of the translation, including accuracy (e.g., mistranslation, omission, addition), fluency (e.g., grammar, spelling, punctuation, inconsistency), locale convention (e.g., format for names, currency, address), terminology (e.g., inappropriate or inconsistent usage), and style (e.g., stylistic problems).
Table 14: Datasets on Machine Translation task.
Dataset Size Aspect
DialSummEval (Gao and Wan, 2022) 5600 Coherence: Measure the quality of all sentences of the summary collectively, to fit together and sound naturally. Consider the quality of the summary as a whole.
Consistency: Measure whether the facts in the summary are consistent with the facts in the dialogue. Consider whether the summary does reproduce all facts accurately and does not make up untrue information.
Fluency: Measure the quality of individual sentences of the summary, whether they are well-written and grammatically correct. Consider the quality of individual sentences.
Relevance: Measure how well the summary captures the key points of the dialogue. Consider whether all and only the important aspects are contained in the summary.
frank (Pagnoni et al., 2021) 2246 Factuality: Measure whether the facts in the summary are correct according to the article.
Newsroom (Grusky et al., 2018) 1680 Coherence: Do phrases and sentences of the summary fit together and make sense collectively?
Fluency: Are the individual sentences of the summary well-written and grammatical?
Informativeness: How well does the summary capture the key points of the article?
Relevance: Are the details provided by the summary consistent with details in the article?
OpenAI (Stiennon et al., 2020) 34197 Accuracy: Does the factual information in the summary accurately match the post? A summary is accurate if it doesn’t say things that aren’t in the post, it doesn’t mix up people, and generally is not misleading.
Coherence: How coherent is the summary on its own? A summary is coherent if, when read by itself, it’s easy to understand and free of English errors. A summary is not coherent if it’s difficult to understand what the summary is trying to say. Generally, it’s more important that the summary is understandable than it being free of grammar errors.
Coverage: How well does the summary cover the important information in the post? A summary has good coverage if it mentions the main information from the post that’s important to understand the situation described in the post. A summary has poor coverage if someone reading only the summary would be missing several important pieces of information about the situation in the post. A summary with good coverage should also match the purpose of the original post (e.g. to ask for advice).
Overall Quality: How good is the summary overall at representing the post? This can encompass coherence (how coherent is the summary on its own), accuracy (does the factual information in the summary accurately match the post), and coverage (how well does the summary cover the important information in the post) of the summary, as well as other important aspects.
QAGS (Wang et al., 2020) 474 Factual Consistency: Is the summary factually consistent with the article? Non-grammatical sentences should be considered not consistent and copies of article sentences should be considered consistent.
Table 15: Datasets on Summarization task. (Part 1)
Dataset Size Aspect
OpinSummEval (Shen and Wan, 2023) 5600 Aspect Relevance: Measure whether the mainly discussed aspects in the reviews are covered exactly by the summary. It focuses on whether the summary correctly reflects the mainly discussed aspects in the reviews.
Readability: Measure whether the summary is fluent and informative. It focuses on whether the summary is well-written and valuable.
Self-coherence: Measure whether the summary is consistent within itself in terms of sentiments and aspects. It focuses on whether the summary is coherent and does not reflect conflicting opinions.
Sentiment Consistency: Measure whether the summary is consistent with the reviews in terms of sentiments for each aspect. It focuses on whether the summary aspect-wisely captures the main sentiment in the reviews.
PolyTope (Huang et al., 2020a) 1268 Overall Quality: The overall quality of the summary, including accuracy and fluency. Accuracy-related issues refer to the extent to which the summary does not match the article, including unnecessary snippets, missing key points, content unfaithful to the article, content not present in the article and factually incorrect, and statements that contradict the article in attitudes (e.g. from positive to negative). Fluency-related issues refer to the linguistic quality of the summary, which is independent of the relationship between the article and the summary, including unnecessary repetition, problems in the word form, and problems in word order.
protagolabs (Sottana et al., 2023) 1600 Coherence: Measure the quality of all sentences of the summary collectively, to fit together and sound naturally. Consider the quality of the summary as a whole.
Consistency: Measure whether the facts in the summary are consistent with the facts in the article. Consider whether the summary does reproduce all facts accurately and does not make up untrue information.
Fluency: Measure the quality of individual sentences of the summary, whether they are well-written and grammatically correct. Consider the quality of individual sentences.
Relevance: Measure how well the summary captures the key points of the article. Consider whether all and only the important aspects are contained in the summary.
SummEval (Fabbri et al., 2021) 6400 Coherence: Measure the quality of all sentences of the summary collectively, to fit together and sound naturally. Consider the quality of the summary as a whole.
Consistency: Measure whether the facts in the summary are consistent with the facts in the article. Consider whether the summary does reproduce all facts accurately and does not make up untrue information.
Fluency: Measure the quality of individual sentences of the summary, whether they are well-written and grammatically correct. Consider the quality of individual sentences.
Relevance: Measure how well the summary captures the key points of the article. Consider whether all and only the important aspects are contained in the summary.
SummEval-OP (Siledar et al., 2024) 2912 Aspect Coverage: The summary should cover all the aspects that are majorly being discussed in the reviews. The summary should be penalized if it misses out on an aspect that was majorly discussed in the reviews and awarded if it covers all.
Coherence: Measure the collective quality of all sentences. The summary should be well-structured and well-organized. The summary should not just be a heap of related information, but should build from sentences to a coherent body of information.
Faithfulness: Every piece of information mentioned in the summary should be verifiable/supported/inferred from the reviews only. The summary should be penalized if any piece of information is not verifiable/supported/inferred from the reviews or if the summary overgeneralizes something.
Fluency: Measure the quality of the summary in terms of grammar, spelling, punctuation, capitalization, word choice, and sentence structure and should contain no errors. The summary should be easy to read, follow, comprehend and should contain no errors.
Relevance: The summary should not contain opinions that are either not consensus or important. The summary should include only important opinions from the reviews. The summary should be penalized if it contains redundancies and unimportant information.
Sentiment Consistency: All the aspects discussed in the summary should accurately reflect the consensus sentiment of the corresponding aspects from the reviews. The summary should be penalized if it does not cover accurately the sentiment regarding any aspect within the summary.
Specificity: The summary should avoid containing generic opinions. All the opinions within the summary should contain detailed and specific information about the consensus opinions. The summary should be penalized for missing out details and should be awarded if they are specific.
Table 16: Datasets on Summarization task. (Part 2)
Dataset Size Aspect
ASSET (Alva-Manchego et al., 2020) 300 Adequacy (or Meaning Preservation): The simplified sentence should adequately express the meaning of the original sentence, perhaps omitting the least important information.
Fluency (or Grammaticality): The simplified sentence should be fluent, and there are no grammatical errors.
Simplicity: The simplified sentence should be easier to understand than the original sentence.
Fusion (Schwarzer et al., 2021) 10490 Adequacy: To which degree does the simplified sentence retain the meaning of the original sentence?
Simplicity: To which degree is the simplified sentence simpler than the original sentence?
HSplit (Sulem et al., 2018c) 7560 Grammaticality: Is the simplified text fluent and grammatical?
Meaning Preservation: Does the simplified text preserve the meaning of the original text?
Simplicity: Is the simplified text simpler than the original text?
Structural Simplicity: Is the simplified text simpler than the original text, ignoring the complexity of the words?
Human-likert (Scialom et al., 2021) 336 Fluency: How fluent is the simplified text?
Meaning Preservation: How well does the simplified text express the original meaning?
Simplicity: To what extent is the simplified text easier to read and understand?
LENS (Maddela et al., 2023) 3840 Overall Quality: The simplified sentence should be fully simplified, entirely fluent, and preserve the core meaning of the original sentence.
Adequacy: The simplified sentence should adequately express the meaning of the original sentence, perhaps omitting the least important information.
Fluency: The simplified sentence should be fluent, and there are no grammatical errors.
Simplicity: The simplified sentence should be easier to understand than the original sentence.
metaeval (Alva-Manchego et al., 2021) 1800 Adequacy (or Meaning Preservation): Judge by looking at both the original and simplified texts, and judge whether or not the changes made preserve the original meaning.
Fluency (or Grammaticality): Judge by looking solely at the simplified text. Mainly consider the grammatical and/or spelling errors, but also ’how well’ (or natural) the text reads. Do not take capitalization into consideration.
Simplicity: Judge by looking at both the original and simplified texts, and judge whether or not the changes made the simplified text easier to understand than the original text.
protagolabs (Sottana et al., 2023) 1200 Fluency (or Grammaticality): Whether the simplified text remains grammatical and understandable?
Semantics (or Adequacy): Whether the meaning is preserved further to the simplification?
Simplicity: Whether the simplified text is simpler than the original text?
SAMSA (Sulem et al., 2018b) 1500 Grammaticality: Is the simplified text grammatical?
Meaning Preservation: Does the simplified text add information or remove important information, compared to the original text?
Structural Simplicity: Is the simplified text simpler than the original text, ignoring the complexity of the words?
Table 17: Datasets on Text Simplification task.
Prompts and Instructions
###Instruction###
Please act as an impartial and helpful evaluator for natural language generation (NLG), and the audience is an expert in the field.
Your task is to evaluate the quality of {task} strictly based on the given evaluation criterion.
Begin the evaluation by providing your analysis concisely and accurately, and then on the next line, start with "Rating:" followed by your rating on a Likert scale from 1 to 5 (higher means better).
You MUST keep to the strict boundaries of the evaluation criterion and focus solely on the issues and errors involved; otherwise, you will be penalized.
Make sure you read and understand these instructions, as well as the following evaluation criterion and example content, carefully.
###Evaluation Criterion###
{aspect}
###Example###
{source_des}:
{source}
{target_des}:
{target}
###Your Evaluation###
Table 18: Prompts and instructions used for evaluating NLG tasks.
Prompts and Instructions
###Instruction###
You are a professional and helpful evaluator for natural language generation (NLG).
You will be given an example of Review Generation task, which includes an aspect description and a review based on it.
Your task is to evaluate the quality of the review strictly based on the given evaluation criterion.
Your evaluation MUST begin with the accurate analysis, followed by ’Rating:’ and then include the corresponding evaluation rating.
Your rating MUST be an integer ranging from 1 to 5, following a five-point Likert scale. (higher means better)
You MUST keep to the strict boundaries of the given evaluation criterion and focus ONLY on the issues and errors involved; otherwise, you will be penalized.
Make sure you read and understand these instructions, as well as the following evaluation criterion and example content, carefully.
###Evaluation Criterion###
Aspect Alignment: Is the review strictly aligned with and solely based on the corresponding aspect description, without mentioning any other points out of scope?
###Example###
Aspect Description:
{aspect}
Review:
{analysis}
###Your Evaluation###
Table 19: Prompts and instructions used for re-evaluating evaluations based on the alignment between evaluation analyses and aspects.
Prompts and Instructions
###Instruction###
You are a professional and helpful evaluator for natural language generation (NLG).
You will be given an example of Review Generation task, which includes a review of the {target_des} and a polarity.
Your task is to evaluate the quality of the review strictly based on the given evaluation criterion.
Your evaluation MUST begin with the accurate analysis, followed by ’Rating:’ and then include the corresponding evaluation rating.
Your rating MUST be an integer ranging from 1 to 5, following a five-point Likert scale. (higher means better)
You MUST keep to the strict boundaries of the given evaluation criterion and focus ONLY on the issues and errors involved; otherwise, you will be penalized.
Make sure you read and understand these instructions, as well as the following evaluation criterion and example content, carefully.
###Evaluation Criterion###
Polarity Consistency: Is the polarity of the review towards the {target_des} consistent with the given polarity (including negative, slightly negative, neutral, slightly positive, and positive)?
###Example###
Polarity:
{rating_to_polarity}
Review:
{analysis}
###Your Evaluation###
Table 20: Prompts and instructions used for re-evaluating evaluations based on the alignment between evaluation analyses and ratings.
Appendix C Training Details
The sizes of training data used in our supervised fine-tuning and preference alignment are 67,180 and 10,000, respectively. In the construction of the latter, we prioritize selecting preference pairs with larger differences in evaluation ratings, as intuitively, their preference relations are more reliable. In addition, following the construction method of the training data, we built an additional validation set based on the NLG Eval corpus, comprising a total of 2,479 samples, to find the most suitable training hyperparameters. During supervised fine-tuning, the learning rate is 1e-5, the batch size is 128, the training epoch is 3, and the optimizer is Adam. And in preference alignment, the learning rate is 3e-6, the batch size is 128, the training epoch is 1, $\alpha$ is 1.0, and the optimizer is Adam.
Appendix D Complete Results of Six NLG Tasks
We display the complete results of our Themis and other models on six NLG tasks respectively in Table 21 to Table 26.
Method Coherence Consistency Fluency Relevance Average
$\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$
Traditional Metrics
BLEU 0.062 0.044 0.048 0.040 0.046 0.036 0.145 0.108 0.075 0.057
ROUGE 0.107 0.080 0.145 0.123 0.113 0.093 0.241 0.183 0.152 0.120
BARTScore 0.474 0.367 0.266 0.220 0.258 0.214 0.318 0.243 0.329 0.261
BERTScore 0.285 0.220 0.151 0.122 0.186 0.154 0.302 0.232 0.231 0.182
BLEURT 0.150 0.112 0.089 0.074 0.133 0.107 0.238 0.178 0.152 0.118
CometKiwi 0.353 0.273 0.151 0.124 0.207 0.170 0.203 0.151 0.228 0.180
UniEval 0.575 0.442 0.446 0.371 0.449 0.371 0.426 0.325 0.474 0.377
Prompt LLM
G-Eval (GPT-3.5) 0.440 0.335 0.386 0.318 0.424 0.347 0.385 0.293 0.409 0.323
G-Eval (GPT-4) 0.582 0.457 0.507 0.425 0.455 0.378 0.548 0.433 0.523 0.423
GPT-3.5 0.459 0.371 0.393 0.331 0.355 0.296 0.455 0.363 0.415 0.340
GPT-4 0.540 0.434 0.531 0.464 0.480 0.409 0.491 0.395 0.511 0.426
AUTOCALIBRATE (Liu et al., 2024b) 0.570 0.493 0.500 0.467 0.487 0.452 0.560 0.483 0.529 0.474
CoAScore(n=10) (Gong and Mao, 2023b) 0.541 0.419 0.339 0.299 0.367 0.308 0.478 0.379 0.431 0.351
HD-EVAL-NN (Liu et al., 2024c) 0.657 - 0.451 - 0.435 - 0.599 - 0.535 -
Fine-tuned LLM
X-Eval 0.530 0.382 0.428 0.340 0.461 0.365 0.500 0.361 0.480 0.362
Prometheus 0.150 0.126 0.150 0.137 0.189 0.168 0.164 0.138 0.163 0.142
AUTO-J 0.245 0.203 0.131 0.121 0.154 0.141 0.262 0.222 0.198 0.172
TIGERScore 0.381 0.318 0.427 0.387 0.363 0.327 0.366 0.304 0.384 0.334
InstructScore 0.328 0.276 0.232 0.213 0.260 0.237 0.211 0.179 0.258 0.226
Themis (ours) 0.566 0.485 0.600 0.566 0.571 0.533 0.474 0.412 0.553 0.499
Table 21: Complete results on SummEval for summarization task.
Method Context Maintenance Interestingness Knowledge Use Naturalness Average
$r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$
Traditional Metrics
BLEU 0.370 0.374 0.406 0.454 0.281 0.369 0.366 0.356 0.356 0.388
ROUGE 0.400 0.376 0.452 0.488 0.339 0.423 0.381 0.360 0.393 0.412
BARTScore 0.119 0.165 0.069 0.059 0.031 0.053 0.050 0.065 0.067 0.086
BERTScore 0.395 0.383 0.439 0.449 0.330 0.378 0.388 0.366 0.388 0.394
BLEURT 0.401 0.408 0.431 0.427 0.321 0.364 0.383 0.354 0.384 0.388
comet22 0.491 0.496 0.544 0.544 0.407 0.450 0.489 0.492 0.483 0.496
CometKiwi 0.334 0.327 0.369 0.355 0.331 0.309 0.380 0.368 0.353 0.340
UniEval 0.595 0.613 0.557 0.605 0.536 0.575 0.444 0.514 0.533 0.577
Prompt LLM
G-Eval (GPT-3.5) 0.519 0.544 0.660 0.691 0.586 0.567 0.532 0.539 0.574 0.585
G-Eval (GPT-4) 0.594 0.605 0.627 0.631 0.531 0.551 0.549 0.565 0.575 0.588
GPT-3.5 0.550 0.531 0.651 0.648 0.653 0.581 0.515 0.550 0.592 0.578
GPT-4 0.680 0.680 0.822 0.779 0.810 0.786 0.769 0.739 0.770 0.746
CoAScore(n=20) 0.539 0.553 0.578 0.595 - - 0.558 0.596 - -
HD-EVAL-NN 0.584 0.607 0.682 0.701 0.549 0.568 0.648 0.674 0.616 0.638
Fine-tuned LLM
X-Eval 0.558 0.622 0.449 0.593 0.734 0.728 0.417 0.478 0.539 0.605
Prometheus 0.451 0.465 0.495 0.473 0.437 0.412 0.355 0.384 0.435 0.434
AUTO-J 0.452 0.449 0.490 0.459 0.339 0.357 0.425 0.437 0.427 0.425
TIGERScore 0.417 0.438 0.328 0.333 0.137 0.138 0.455 0.477 0.334 0.346
InstructScore 0.299 0.297 0.264 0.233 0.140 0.102 0.374 0.332 0.269 0.241
Themis (ours) 0.639 0.644 0.790 0.766 0.778 0.761 0.727 0.729 0.733 0.725
Table 22: Complete results on Topical-Chat for dialogue response generation task.
Method SFHOT INF. SFHOT NAT. SFRES INFO. SFRES NAT. Average
$\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$
Traditional Metrics
BLEU 0.070 0.054 0.055 0.040 -0.023 -0.018 -0.004 -0.004 0.024 0.018
ROUGE 0.107 0.082 0.075 0.055 0.118 0.090 0.105 0.078 0.101 0.076
BARTScore 0.211 0.162 0.130 0.094 0.265 0.201 0.226 0.165 0.208 0.156
BERTScore 0.135 0.104 0.126 0.093 0.157 0.120 0.139 0.102 0.139 0.105
BLEURT 0.219 0.171 0.229 0.171 0.244 0.186 0.282 0.211 0.244 0.184
CometKiwi 0.220 0.169 0.235 0.172 0.203 0.153 0.345 0.252 0.251 0.186
UniEval 0.249 0.191 0.320 0.238 0.225 0.169 0.333 0.247 0.282 0.211
Prompt LLM
GPT-3.5 0.242 0.196 0.294 0.220 0.304 0.250 0.385 0.291 0.306 0.239
GPT-4 0.302 0.263 0.359 0.283 0.213 0.178 0.405 0.316 0.320 0.260
AUTOCALIBRATE 0.357 0.313 0.440 0.383 0.315 0.272 0.416 0.351 0.382 0.330
Fine-tuned LLM
Prometheus 0.169 0.141 0.211 0.171 0.161 0.134 0.150 0.122 0.173 0.142
Auto-J 0.176 0.152 0.127 0.106 0.179 0.153 0.084 0.070 0.141 0.120
TIGERScore 0.215 0.191 0.204 0.175 0.160 0.141 0.221 0.191 0.200 0.175
InstructScore 0.222 0.194 0.273 0.231 0.194 0.164 0.300 0.251 0.247 0.210
Themis (ours) 0.259 0.226 0.380 0.321 0.298 0.258 0.395 0.332 0.333 0.284
Table 23: Complete results on SFRES & SFHOT for Data to Text task.
Method CNN-DM XSUM Average
$r$ $\rho$ $\tau$ $r$ $\rho$ $\tau$ $r$ $\rho$ $\tau$
Traditional Metrics
BARTScore 0.732 0.680 0.555 0.175 0.171 0.139 0.454 0.425 0.347
CometKiwi 0.176 0.158 0.123 0.027 0.030 0.025 0.101 0.094 0.074
UniEval 0.682 0.662 0.532 0.461 0.488 0.399 0.572 0.575 0.466
Prompt LLM
G-Eval (GPT-4) 0.477 0.516 0.410 0.211 0.406 0.343 0.344 0.461 0.377
G-Eval (GPT-3.5) 0.631 0.685 0.591 0.558 0.537 0.472 0.595 0.611 0.532
GPT-3.5 0.454 0.514 0.417 0.279 0.348 0.295 0.366 0.431 0.356
GPT-4 0.735 0.746 0.626 0.541 0.528 0.439 0.638 0.637 0.532
AUTOCALIBRATE 0.740 0.744 0.663 0.662 0.662 0.662 0.701 0.703 0.663
Fine-tuned LLM
Auto-J 0.291 0.238 0.214 0.225 0.214 0.203 0.258 0.226 0.209
TIGERScore 0.574 0.562 0.479 0.424 0.445 0.412 0.499 0.504 0.446
InstructScore 0.287 0.278 0.233 -0.096 -0.134 -0.119 0.095 0.072 0.057
Themis (ours) 0.747 0.761 0.680 0.599 0.607 0.546 0.673 0.684 0.613
Table 24: Complete results on QAGS for factuality evaluation task.
Method ROC WP Average
$r$ $\rho$ $\tau$ $r$ $\rho$ $\tau$ $r$ $\rho$ $\tau$
Traditional Metrics
BLEU 0.034 0.035 0.022 0.009 0.029 0.021 0.021 0.032 0.009
ROUGE 0.012 0.000 0.006 0.003 -0.004 -0.004 0.008 -0.002 0.156
BARTScore 0.344 0.330 0.273 0.403 0.370 0.306 0.373 0.350 0.260
BERTScore 0.288 0.269 0.222 0.293 0.300 0.248 0.291 0.285 0.163
BLEURT 0.189 0.155 0.123 0.144 0.122 0.104 0.166 0.138 0.221
CometKiwi 0.246 0.218 0.176 0.289 0.283 0.238 0.268 0.251 0.176
Prompt LLM
GPT-3.5 0.372 0.363 0.312 0.312 0.294 0.258 0.342 0.328 0.295
GPT-4 0.590 0.578 0.518 0.382 0.368 0.336 0.486 0.473 0.260
CoAScore(n=20) - - - - - - 0.414 0.411 0.315
Fine-tuned LLM
Prometheus 0.031 0.013 0.014 0.001 0.001 0.003 0.016 0.007 0.146
Auto-J 0.460 0.454 0.410 0.308 0.306 0.278 0.384 0.380 0.284
TIGERScore 0.283 0.271 0.221 0.196 0.191 0.158 0.240 0.231 0.207
InstructScore 0.383 0.368 0.316 0.258 0.228 0.194 0.321 0.298 0.168
Themis (ours) 0.637 0.607 0.551 0.507 0.495 0.452 0.572 0.551 0.501
Table 25: Complete results on MANS for story generation task.
Method WMT23
$r$ $\rho$ $\tau$
Traditional Metrics
BLEU -0.130 0.021 0.018
ROUGE 0.081 0.151 0.117
BARTScore 0.091 0.118 0.093
BERTScore 0.123 0.219 0.170
BLEURT 0.163 0.263 0.208
CometKiwi 0.413 0.343 0.273
Prompt LLM
GPT-3.5 0.388 0.347 0.278
GPT-4 0.496 0.437 0.361
Fine-tuned LLM
Prometheus 0.144 0.129 0.107
Auto-J 0.128 0.104 0.087
TIGERScore 0.277 0.248 0.211
InstructScore 0.213 0.219 0.181
Themis (ours) 0.431 0.405 0.357
Table 26: Complete results on WMT23 zh-en for machine translation task.